OpenStudy (anonymous):

find dy/dx explicitly as dy/dx=g(x,y) if x^3+y^2+4y=4

OpenStudy (anonymous):

\[3x^2\div(-2y-4)=(dy \div dx)\]

OpenStudy (ash2326):

we have \(x^3+y^2+4y=4\) differentiate with respect to x \( 3x^2+2y dy/dx+4 dy/dx=0\) \((2y+4)dy/dx=-3x^2\) \[dy/dx= -3x^2/(2y+4)\]

OpenStudy (anonymous):

thax loads for your help